[Hallc_running] [New Logentry] Hall C Update (Dec. 8)

gaskelld at jlab.org gaskelld at jlab.org
Fri Dec 8 09:20:02 EST 2017

Logentry Text:
Some issues with the 5th pass RF separator were addressed yesterday, and setup to hall D and B resumed last night. Unfortunately, some new issues with the separator appeared, so the decision was made to setup Hall B and C on owl shift last night.

The Hall C beam was propagating through the hall C arc last night, but could not get past 3C14 (upstream of the green shield wall). There is the suspicion that there might be an obstruction - this is to be determined.

Today, there will be work to assess the RF separator issues and determine the source of the Hall C beam transport issues.

The expectation is that the beam transport issue to hall C will be resolved and setup of Hall C will proceed later today. All shifts should report as scheduled.  


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3501142
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20171208/fd336c06/attachment.html>

More information about the Hallc_running mailing list